integration of (1/e power x).dxplease answer this question as soon as possible.

Grade:12

1 Answers

Satarup Dutta
6 years ago
Take the integral: integral e^(-x) dxFor the integrand e^(-x), substitute u = -x and du = - dx: = - integral e^u duThe integral of e^u is e^u: = -e^u + constantSubstitute back for u = -x:Answer: | | = -e^(-x) + constant